OpenStudy (anonymous):

solve this system of linear equations separate the x- and y-values with a comma -7x+6y=-5 2x+4y=-10

OpenStudy (amistre64):

i wonder if we can multiply the top by 2 and solve that way

OpenStudy (anonymous):

that sounds....plausible. lets try it!

OpenStudy (amistre64):

-7x+6y=-5 ; *2 2x+4y=-10 -14x+12y=-10 2x+ 4y=-10 -------------- -12x+16y= 0 ; prolly not

OpenStudy (amistre64):

doh!! forgot to do real math on it

OpenStudy (anonymous):

woah buddy, this is no place for real math!

OpenStudy (amistre64):

-7x+6y=-5 ; *2 2x+4y=-10 -14x+12y=-10 2x+ 4y=-10 -14x+12y= 2x+4y -14x-2x+12y-4y= 0 -16x +8y = 0 8y = 16x y = 2x ; if anything y might be twice x

OpenStudy (amistre64):

-7x+6y=-5 ; *2/7 2x+4y=-10 -2x+12y/7 =-10/7 2x+4y=-10 -2x+12y/7 =-10/7 2x+28y/7 =-70/7 ------------------ 40y/7 = -80/7 y = -2 ?

OpenStudy (amistre64):

which may mean that x = -1 -7(-1)+6(-2)=-5 2(-1)+4(-2)=-10 7 -12 =-5 -2 -8=-10 lol .. it does

OpenStudy (anonymous):

yea it wuz that da answer was -1,-2

OpenStudy (anonymous):

thanks